# TRIDENT: Tri-Modal Molecular Representation Learning with Taxonomic Annotations and Local Correspondence

### v1

@@ Abstract
Molecular property prediction aims to learn representations that map chemical structures to functional properties. While multimodal learning has emerged as a powerful paradigm to learn molecular representations, prior works have largely overlooked textual and taxonomic information of molecules for representation learning. We introduce TRIDENT, a novel framework that integrates molecular SMILES, textual descriptions, and taxonomic functional annotations to learn rich molecular representations. To achieve this, we curate a comprehensive dataset of molecule-text pairs with structured, multi-level functional annotations. Instead of relying on conventional contrastive loss, TRIDENT employs a volume-based alignment objective to jointly align tri-modal features at the global level, enabling soft, geometry-aware alignment across modalities. Additionally, TRIDENT introduces a novel local alignment objective that captures detailed relationships between molecular substructures and their corresponding sub-textual descriptions. A momentum-based mechanism dynamically balances global and local alignment, enabling the model to learn both broad functional semantics and fine-grained structure-function mappings. TRIDENT achieves state-of-the-art performance on 11 downstream tasks, demonstrating the value of combining SMILES, textual, and taxonomic functional annotations for molecular property prediction.
@@ 1 Introduction
Molecular representation learning, which converts complex chemical structures into computational features, has been instrumental in advancing various aspects of drug discovery including virtual screening, and molecular design (liu2022molecular, ; chibani2020machine, ; shen2019molecular, ). Multi-modal molecular models further enhance representation quality by integrating structural, textual, and functional information, enabling better generalization and predictive performance (liu2023multi, ). These approaches hold promise for unlocking deeper insights into chemical space and accelerating the discovery of therapeutic compounds with desired properties.
@@
However, current multimodal approaches (luo2023molfm, ; su2022molecular, ) face three key limitations: (1) Overlooking fine-grained annotations across taxonomies: Most existing methods simplify the representation of molecules by focusing on unified functional descriptions, neglecting the nuanced annotations provided by different taxonomic systems. As illustrated in Figure 2, the same molecule may have distinct emphases depending on the taxonomy: for example, the LOTUS Tree (rutz2022lotus, ) taxonomy highlights natural product classifications, whereas the MeSH (Medical Subject Headings) Tree (lipscomb2000medical, ) taxonomy emphasizes medical functionalities of the same molecule. Ignoring these taxonomy-specific, fine-grained annotations risks reducing molecules to flat entities, thereby failing to capture the multi-faceted and structured nature of chemical functions. (2) Alignment limitations: Aligning modalities such as molecular structures, textual descriptions, and taxonomic functional annotations is inherently complex. Existing methods rely on pairwise alignment schemes anchored to a single modality, which struggle to model the interdependencies across all modalities (zhu2023languagebind, ; liu2022umt, ; xu2023mplug, ; chen2023vast, ), particularly when one modality encodes nested or multi-level information (cicchetti2024gramian, ). (3) Neglect of local correspondences: Many approaches focus exclusively on molecule-level alignment, disregarding the fine-grained relationships between molecular substructures (e.g., functional groups) and their corresponding sub-textual descriptions. This omission limits the expressivity of the learned representations and constrains their applicability in molecular property prediction tasks.
@@
To address these limitations, we introduce the TRIDENT (Tri-modal Representation Integrating Descriptions, Entities, and Taxonomies) framework for molecules that jointly models molecular SMILES, textual descriptions, and multi-faceted Hierarchical Taxonomic Annotation (HTA). Central to TRIDENT is the HTA modality, which organizes molecular function across hierarchical classification levels. We curate a high quality dataset of 47,269 <SMILES, Text, HTA> triplets from PubChem (kim2016pubchem, ), annotated under 32 classification systems. To tackle the challenge of aligning these diverse modalities, TRIDENT leverages a volume-based contrastive loss, enabling soft, geometry-aware alignment of all three modalities. While recently proposed for general-purpose modality alignment (cicchetti2024gramian, ), we extend this formulation to the molecular domain for the first time, where the modalities are structurally diverse and include taxonomic semantic labels. Furthermore, TRIDENT introduces a novel local alignment module that links molecular substructures to their associated sub-textual descriptions, capturing fine-grained structure–function relationships. A momentum-based balancing mechanism dynamically integrates global and local alignments to optimize the representation learning process (see Figure 1 for an overview).
@@ 2.1 Molecule-Text Multimodal Learning
Recent advancements in molecular representation learning have demonstrated the power of multimodal approaches that integrate information from molecular graphs, SMILES strings, and textual descriptions to enhance property prediction and drug discovery. Graph Neural Networks (GNNs) have become the backbone of graph-based methods (an2024causal, ), with models like GROVER (rong2020self, ) and MolCLR wang2022molecular leveraging contrastive learning to produce richer molecular embeddings. Multimodal models such as KV-PLM (zeng2022deep, ) and MolT5 (edwards2022translation, ) treat SMILES and text as separate languages for pre-training via auto-encoding objectives, while MoMu (su2022molecular, ) and MoleculeSTM (liu2023multi, ) utilize independent encoders with cross-modal contrastive learning to align graphs and texts. MolFM (luo2023molfm, ) extends this paradigm by incorporating molecular structures, biomedical texts, and knowledge graphs to capture more comprehensive molecular relationships. However, despite this progress, the textual modality in existing models often derives from unstructured or single-layered descriptions, limiting the capacity to represent molecular functions across diverse biological roles and hierarchical categories. This lack of structured semantic alignment limits the ability of models to reason over complex molecular behaviors and relationships. Our work addresses this gap by introducing a high quality dataset to incorporate hierarchical taxonomic annotations for molecules, learning fine-grained hierarchical molecule-function relationships.
@@ 2.2 Contrastive Learning for Multimodal Alignment
Contrastive learning has emerged as a powerful strategy for aligning representations across modalities. Seminal models such as CLIP (radford2021learning, ) demonstrated effective image-text alignment, inspiring extensions to other domains including audio (CLAP) (elizalde2023clap, ), video (CLIP4Clip) (luo2022clip4clip, ), and point clouds (PointCLIP) (zhang2022pointclip, ). These models typically learn by pulling semantically similar cross-modal pairs closer while pushing dissimilar ones apart (zhong2024compositional, ). More recent approaches such as CLIP4VLA (ruan2023accommodating, ), ImageBind (girdhar2023imagebind, ), and LanguageBind (zhu2023languagebind, ) explore multimodal fusion, often anchoring learning around a central modality like images or text. GRAM (cicchetti2024gramian, ) advances this direction by introducing geometry-aware volume based contrastive objective, but it primarily focuses on audio-video-text pairs without structured semantic hierarchies. In the field of bioinformatics, modality fusion has also shown significant potential, being applied to tasks such as protein structure prediction and drug discovery (jiang2024gte, ; jiang2024alphaepi, ). Unlike existing methods, our TRIDENT framework tackles the unique challenges of molecule-text alignment by incorporating hierarchical taxonomic relationships to capture functional semantics, and introducing global and local alignment modules with momentum-based mechanism. This enables fine-grained substructure-function correspondence and a richer multimodal embedding space tailored to molecular understanding.
@@ 3.1 Hierarchical Taxonomic Annotation (HTA)
To enable structured, hierarchical molecular representations, we introduce the Hierarchical Taxonomic Annotation (HTA) framework, which organizes molecular functions across multiple classification levels. This setup allows the model to capture fine-grained, hierarchical semantics essential for understanding complex molecular properties and their biological roles. We curate a high-quality dataset of 47,269 <SMILES, Text, HTA> triplets sourced from PubChem (kim2016pubchem, ). As shown in Figure  2, these triplets are annotated across 32 diverse hierarchical classification systems, providing a comprehensive, multi-level understanding of molecular behavior. Figure 2 illustrates the construction pipeline for HTA. Beginning with a molecule’s SMILES representation, the molecule is queried against PubChem (kim2016pubchem, ). This yields a set of traditional functional descriptions, which are typically concise, ontology-aware summaries based on cheminformatics rules. For example, citronellol is described as a monoterpenoid… with a role as a plant metabolite. While such descriptors are chemically accurate, they often lack broader context, such as ecological origin, industrial relevance, or toxicological implications.
To address this limitation, we augment the molecule’s annotation space through structured taxonomic enrichment by mapping it into multiple biological and chemical taxonomies. For example, the LOTUS Tree (rutz2022lotus, ) highlights natural product classifications, whereas the MeSH (Medical Subject Headings) Tree (lipscomb2000medical, ) emphasizes medical functionalities of the same molecule. Through this multi-perspective approach, these hierarchies expand the molecular profile beyond flat descriptors into deeply nested semantic trees spanning chemistry, biology, and pharmacology.
In the final stage, we leverage a GPT-4o (achiam2023gpt, ; ouyang2022training, ) to synthesize the retrieved structured annotations into a high-fidelity, human-readable HTA. Unlike traditional descriptors, HTAs encode multi-perspective knowledge: they trace the chemical derivation (e.g., from citronellal), mention natural sources (e.g., rose oil), functional applications (e.g., fragrance in various products), and regulatory or biomedical associations (e.g., environmental protection agencies, blood exposome). This generative synthesis is guided by structural prompts and validated by domain experts to ensure factual accuracy and interoperability.
@@ 3.2 Geometry-based Global Alignment
Traditional multimodal approaches typically rely on pairwise similarity metrics such as cosine similarity: $\text{cos}(\theta_{ij})=\frac{\langle M_{i},M_{j}\rangle}{\|M_{i}\|\cdot\|M_{j}\|}.$ However, these methods often anchor one modality and align others to it independently, failing to capture higher-order relationships across all modalities. To address this, GRAM (cicchetti2024gramian, ) introduced a geometry-based alignment approach that uses the volume of the parallelotope spanned by modality vectors as a global measure of alignment. Specifically, for three normalized embeddings $m$, $t$, and $h$, the volume of the parallelotope is computed as $\text{Vol}(m,t,h)=\sqrt{1-\langle m,t\rangle^{2}-\langle m,h\rangle^{2}-\langle t,h\rangle^{2}+2\langle m,t\rangle\langle t,h\rangle\langle h,m\rangle},$ which reflects the overall geometric alignment of the embeddings. The volume shrinks as the modalities converge and grows as they diverge. Unlike pairwise contrastive learning methods, this formulation was shown to capture the global structure of cross-modal interactions in a principled and scalable way for audio-video-text pairs (cicchetti2024gramian, ).
@@
Following the approach introduced in GRAM (cicchetti2024gramian, ), we construct a global contrastive objective over the three modalities—SMILES, traditional text descriptions, and HTA annotations. Each modality is processed through a modality-specific encoder followed by a modality-specific projection head (implemented as a three-layer MLP) to map the embeddings into a shared latent space, yielding embeddings $m$, $t$, and $h$ for SMILES, text, and HTA, respectively.
@@ 3.3 Fine-grained Local Alignment
To address this limitation, we introduce a local alignment contrastive loss that complements the global volume-based objective. Unlike GRAM (cicchetti2024gramian, ), which operates solely at the level of full modality embeddings, our method leverages the compositional nature of molecules to align substructures with their semantic counterparts in text and taxonomy.
@@ 3.3.1 Functional Group-Level Representation
To enable fine-grained alignment, we construct a high-quality dataset that links functional group structures with their corresponding semantic descriptions. Using RDKit, we screen and identify functionally significant groups frequently found in drug-like molecules. These groups include moieties such as hydroxyls, amines, carboxyls, and aromatic systems. For each group, comprehensive textual descriptions are curated through a hybrid process involving GPT-4o (achiam2023gpt, ) and expert review by professional chemists. This ensures both semantic richness and domain accuracy, resulting in a curated dataset of 85 functional groups paired with high-quality textual annotations.
@@ 4 Experiments
The pre-training dataset is sourced from PubChem, initially following the method described in liu2023multi to obtain approximately 380k SMILES-text pairs. After a series of filtering steps (for detailed processing steps, please refer to the Appendix A) and obtaining HTA information for each molecule, our final pre-training dataset consists of 47,269 SMILES-Text-HTA triplets. The annotations cover various biological roles, molecular functions, mechanisms of action, and multi-level bioactivity information.
@@
We evaluate our model on a broad range of molecular property prediction tasks drawn from two major benchmarks: MoleculeNet (wu2018moleculenet, ) and the Therapeutics Data Commons (TDC) (huang2021therapeutics, ). For MoleculeNet, we include toxicity prediction tasks (BBBP, Tox21, ToxCast), side-effect and clinical toxicity tasks (Sider, ClinTox), and bioactivity classification (MUV, HIV, Bace) tasks. All tasks are framed as binary classification problems, and performance is reported using the ROC-AUC metric.
For the TDC benchmark, we evaluate on DILI, Carcinogens (Languin), and Skin Reaction datasets, reporting both AUC and accuracy following TDC guidelines. Following standard practice, we adopt the scaffold split throughout our methodology to evaluate generalization to novel chemical scaffolds. Each experiment is repeated across three random seeds, and we report the mean and standard deviation. A detailed dataset description can be found in the Appendix C.
@@
We compare our TRIDENT approach against a range of recent state-of-the-art baselines. These include MOLFORMER (ross2022large, ), MegaMolBART (reidenbach2022improving, ) a transformer-based model that uses SMILES representations only, as well as recent multimodal approaches incorporating additional textual and molecular information such as MolFM (luo2023molfm, ), MoMu (su2022molecular, ), MoleculeSTM (liu2023multi, ), and MolCA-SMILES (liu2023molca, ), KV-PLM (zeng2022deep, ) and Atomas (zhang2025atomas, ). A detailed baseline introduction can be found in the Appendix D.

### v2

@@ Abstract
Molecular property prediction aims to learn representations that map chemical structures to functional properties. While multimodal learning has emerged as a powerful paradigm to learn molecular representations, prior works have largely overlooked textual and taxonomic information of molecules for representation learning. We introduce TRIDENT, a novel framework that integrates molecular SMILES, textual descriptions, and taxonomic functional annotations to learn rich molecular representations. To achieve this, we curate a comprehensive dataset of molecule-text pairs with structured, multi-level functional annotations. Instead of relying on conventional contrastive loss, TRIDENT employs a volume-based alignment objective to jointly align tri-modal features at the global level, enabling soft, geometry-aware alignment across modalities. Additionally, TRIDENT introduces a novel local alignment objective that captures detailed relationships between molecular substructures and their corresponding sub-textual descriptions. A momentum-based mechanism dynamically balances global and local alignment, enabling the model to learn both broad functional semantics and fine-grained structure-function mappings. TRIDENT achieves state-of-the-art performance on 18 downstream tasks, demonstrating the value of combining SMILES, textual, and taxonomic functional annotations for molecular property prediction. Our code and data are available at https://github.com/uta-smile/TRIDENT.
@@ 1 Introduction
Molecular representation learning, which converts complex chemical structures into computational features, has been instrumental in advancing various aspects of drug discovery including virtual screening, and molecular design (Liu et al., 2022a; Chibani and Coudert, 2020; Shen and Nicolaou, 2019). Multi-modal molecular models further enhance representation quality by integrating structural, textual, and functional information, enabling better generalization and predictive performance (Liu et al., 2023a). These approaches hold promise for unlocking deeper insights into chemical space and accelerating the discovery of therapeutic compounds with desired properties.
@@
However, current multimodal approaches (Luo et al., 2023; Su et al., 2022) face three key limitations: (1) Overlooking fine-grained annotations across taxonomies: Most existing methods simplify the representation of molecules by focusing on unified functional descriptions, neglecting the nuanced annotations provided by different taxonomic systems. As illustrated in Figure 2, the same molecule may have distinct emphases depending on the taxonomy: for example, the LOTUS Tree (Rutz et al., 2022) taxonomy highlights natural product classifications, whereas the MeSH (Medical Subject Headings) Tree (Lipscomb, 2000) taxonomy emphasizes medical functionalities of the same molecule. Ignoring these taxonomy-specific, fine-grained annotations risks reducing molecules to flat entities, thereby failing to capture the multi-faceted and structured nature of chemical functions. (2) Alignment limitations: Aligning modalities such as molecular structures, textual descriptions, and taxonomic functional annotations is inherently complex. Existing methods rely on pairwise alignment schemes anchored to a single modality, which struggle to model the interdependencies across all modalities (Zhu et al., 2023; Liu et al., 2022b; Xu et al., 2023; Chen et al., 2023), particularly when one modality encodes nested or multi-level information (Cicchetti et al., 2024). (3) Neglect of local correspondences: Many approaches focus exclusively on molecule-level alignment, disregarding the fine-grained relationships between molecular substructures (e.g., functional groups) and their corresponding sub-textual descriptions. This omission limits the expressivity of the learned representations and constrains their applicability in molecular property prediction tasks.
@@
To address these limitations, we introduce the TRIDENT (Tri-modal Representation Integrating Descriptions, Entities, and Taxonomies) framework for molecules that jointly models molecular SMILES, textual descriptions, and multi-faceted Hierarchical Taxonomic Annotation (HTA). Central to TRIDENT is the HTA modality, which organizes molecular function across hierarchical classification levels. We curate a high quality dataset of 47,269 <SMILES, Text, HTA> triplets from PubChem (Kim et al., 2016), annotated under 32 classification systems. To tackle the challenge of aligning these diverse modalities, TRIDENT leverages a volume-based contrastive loss, enabling soft, geometry-aware alignment of all three modalities. While recently proposed for general-purpose modality alignment (Cicchetti et al., 2024), we extend this formulation to the molecular domain for the first time, where the modalities are structurally diverse and include taxonomic semantic labels. Furthermore, TRIDENT introduces a novel local alignment module that links molecular substructures to their associated sub-textual descriptions, capturing fine-grained structure–function relationships. A momentum-based balancing mechanism dynamically integrates global and local alignments to optimize the representation learning process (see Figure 1 for an overview).
@@ 2.1 Molecule-Text Multimodal Learning
Recent advancements in molecular representation learning have demonstrated the power of multimodal approaches that integrate information from molecular graphs, SMILES strings, and textual descriptions to enhance property prediction and drug discovery. Graph Neural Networks (GNNs) have become the backbone of graph-based methods, with models like GROVER (Rong et al., 2020) and MolCLR Wang et al. (2022) leveraging contrastive learning to produce richer molecular embeddings. Multimodal models such as KV-PLM (Zeng et al., 2022) and MolT5 (Edwards et al., 2022) treat SMILES and text as separate languages for pre-training via auto-encoding objectives, while MoMu (Su et al., 2022) and MoleculeSTM (Liu et al., 2023a) utilize independent encoders with cross-modal contrastive learning to align graphs and texts. MolFM (Luo et al., 2023) extends this paradigm by incorporating molecular structures, biomedical texts, and knowledge graphs to capture more comprehensive molecular relationships. However, despite this progress, the textual modality in existing models often derives from unstructured or single-layered descriptions, limiting the capacity to represent molecular functions across diverse biological roles and hierarchical categories. This lack of structured semantic alignment limits the ability of models to reason over complex molecular behaviors and relationships. Our work addresses this gap by introducing a high quality dataset to incorporate hierarchical taxonomic annotations for molecules, learning fine-grained hierarchical molecule-function relationships.
@@ 2.2 Contrastive Learning for Multimodal Alignment
Contrastive learning has emerged as a powerful strategy for aligning representations across modalities. Seminal models such as CLIP (Radford et al., 2021) demonstrated effective image-text alignment, inspiring extensions to other domains including audio (CLAP) (Elizalde et al., 2023), video (CLIP4Clip) (Luo et al., 2022), and point clouds (PointCLIP) (Zhang et al., 2022). These models typically learn by pulling semantically similar cross-modal pairs closer while pushing dissimilar ones apart. More recent approaches such as CLIP4VLA (Ruan et al., 2023), ImageBind (Girdhar et al., 2023), and LanguageBind (Zhu et al., 2023) explore multimodal fusion, often anchoring learning around a central modality like images or text. GRAM (Cicchetti et al., 2024) advances this direction by introducing geometry-aware volume based contrastive objective, but it primarily focuses on audio-video-text pairs without structured semantic hierarchies. In bioinformatics, multimodal learning has shown promise in integrating diverse biological data sources (Jiang et al., 2024a; Dang et al., 2025; Miao et al., 2025), such as molecular structures, protein sequences, and biomedical text, to enhance understanding of complex biological systems and accelerate drug discovery (Jiang et al., 2024b; Ma et al., 2022, 2024). Unlike existing methods, our TRIDENT framework tackles the unique challenges of molecule-text alignment by incorporating hierarchical taxonomic relationships to capture functional semantics, and introducing global and local alignment modules with momentum-based mechanism. This enables fine-grained substructure-function correspondence and a richer multimodal embedding space tailored to molecular understanding.
@@ 3.1 Hierarchical Taxonomic Annotation (HTA)
To enable structured, hierarchical molecular representations, we introduce the Hierarchical Taxonomic Annotation (HTA) framework, which organizes molecular functions across multiple classification levels. This setup allows the model to capture fine-grained, hierarchical semantics essential for understanding complex molecular properties and their biological roles. We curate a high-quality dataset of 47,269 <SMILES, Text, HTA> triplets sourced from PubChem (Kim et al., 2016). As shown in Figure  2, these triplets are annotated across 32 diverse hierarchical classification systems, providing a comprehensive, multi-level understanding of molecular behavior. Figure 2 illustrates the construction pipeline for HTA. Beginning with a molecule’s SMILES representation, the molecule is queried against PubChem (Kim et al., 2016). This yields a set of traditional functional descriptions, which are typically concise, ontology-aware summaries based on cheminformatics rules. For example, citronellol is described as a monoterpenoid… with a role as a plant metabolite. While such descriptors are chemically accurate, they often lack broader context, such as ecological origin, industrial relevance, or toxicological implications.
To address this limitation, we augment the molecule’s annotation space through structured taxonomic enrichment by mapping it into multiple biological and chemical taxonomies. For example, the LOTUS Tree (Rutz et al., 2022) highlights natural product classifications, whereas the MeSH (Medical Subject Headings) Tree (Lipscomb, 2000) emphasizes medical functionalities of the same molecule. Through this multi-perspective approach, these hierarchies expand the molecular profile beyond flat descriptors into deeply nested semantic trees spanning chemistry, biology, and pharmacology.
In the final stage, we leverage a GPT-4o (Achiam et al., 2023; Ouyang et al., 2022) to synthesize the retrieved structured annotations into a high-fidelity, human-readable HTA. Unlike traditional descriptors, HTAs encode multi-perspective knowledge: they trace the chemical derivation (e.g., from citronellal), mention natural sources (e.g., rose oil), functional applications (e.g., fragrance in various products), and regulatory or biomedical associations (e.g., environmental protection agencies, blood exposome). This generative synthesis is guided by structural prompts and validated by domain experts to ensure factual accuracy and interoperability.
@@ 3.2 Geometry-based Global Alignment
Traditional multimodal approaches typically rely on pairwise similarity metrics such as cosine similarity: $\text{cos}(\theta_{ij})=\frac{\langle M_{i},M_{j}\rangle}{\|M_{i}\|\cdot\|M_{j}\|}.$ However, these methods often anchor one modality and align others to it independently, failing to capture higher-order relationships across all modalities. To address this, GRAM (Cicchetti et al., 2024) introduced a geometry-based alignment approach that uses the volume of the parallelotope spanned by modality vectors as a global measure of alignment. Specifically, for three normalized embeddings $m$, $t$, and $h$, the volume of the parallelotope is computed as $\text{Vol}(m,t,h)=\sqrt{1-\langle m,t\rangle^{2}-\langle m,h\rangle^{2}-\langle t,h\rangle^{2}+2\langle m,t\rangle\langle t,h\rangle\langle h,m\rangle},$ which reflects the overall geometric alignment of the embeddings. The volume shrinks as the modalities converge and grows as they diverge. Unlike pairwise contrastive learning methods, this formulation was shown to capture the global structure of cross-modal interactions in a principled and scalable way for audio-video-text pairs (Cicchetti et al., 2024).
@@
Following the approach introduced in GRAM (Cicchetti et al., 2024), we construct a global contrastive objective over the three modalities—SMILES, traditional text descriptions, and HTA annotations. Each modality is processed through a modality-specific encoder followed by a modality-specific projection head (implemented as a three-layer MLP) to map the embeddings into a shared latent space, yielding embeddings $m$, $t$, and $h$ for SMILES, text, and HTA, respectively.
@@ 3.3 Fine-grained Local Alignment
To address this limitation, we introduce a local alignment contrastive loss that complements the global volume-based objective. Unlike GRAM (Cicchetti et al., 2024), which operates solely at the level of full modality embeddings, our method leverages the compositional nature of molecules to align substructures with their semantic counterparts in text and taxonomy.
@@ 3.3.1 Functional Group-Level Representation
To enable fine-grained alignment, we construct a high-quality dataset that links functional group structures with their corresponding semantic descriptions. Using RDKit, we screen and identify functionally significant groups frequently found in drug-like molecules. These groups include moieties such as hydroxyls, amines, carboxyls, and aromatic systems. For each group, comprehensive textual descriptions are curated through a hybrid process involving GPT-4o (Achiam et al., 2023) and expert review by professional chemists. This ensures both semantic richness and domain accuracy, resulting in a curated dataset of 85 functional groups paired with high-quality textual annotations.
@@ 4 Experiments
The pre-training dataset is sourced from PubChem, initially following the method described in Liu et al. (2023a) to obtain approximately 380k SMILES-text pairs. After a series of filtering steps (for detailed processing steps, please refer to the Appendix A) and obtaining HTA information for each molecule, our final pre-training dataset consists of 47,269 SMILES-Text-HTA triplets. The annotations cover various biological roles, molecular functions, mechanisms of action, and multi-level bioactivity information.
@@
We evaluate our model on a broad range of molecular property prediction tasks drawn from two major benchmarks: MoleculeNet (Wu et al., 2018) and the Therapeutics Data Commons (TDC) (Huang et al., 2021). For MoleculeNet, we include 8 classification datasets and 3 regression datasets. The classification tasks comprise toxicity prediction (BBBP, Tox21, ToxCast), side-effect and clinical toxicity prediction (Sider, ClinTox), and bioactivity classification (MUV, HIV, Bace), with performance reported using the ROC-AUC metric. The regression tasks include molecular solubility (ESOL), solvation free energy (FreeSolv), and lipophilicity (Lipophilicity), with performance reported using RMSE.
For the TDC benchmark, we evaluate on 7 datasets, including 6 classification datasets (DILI, Carcinogens (Languin), Skin Reaction, hERG, AMES, and CYP P450 2C19) and 1 regression dataset (Caco-2). For classification tasks, we report both AUC and accuracy following TDC guidelines, while for the regression task, we report RMSE. Following standard practice, we adopt the scaffold split throughout our methodology to evaluate generalization to novel chemical scaffolds. Each experiment is repeated across three random seeds, and we report the mean and standard deviation. A detailed dataset description can be found in Appendix C.
@@
We compare our TRIDENT approach against a range of recent state-of-the-art baselines. These include transformer-based models that use SMILES representations, such as MOLFORMER (Ross et al., 2022), MegaMolBART (Reidenbach et al., 2022), MolXPT (Liu et al., 2023b), BioT5 (Pei et al., 2023), and BioT5+(Pei et al., 2024). We also compare against multimodal approaches incorporating additional textual and molecular information, including MolFM(Luo et al., 2023), MoMu (Su et al., 2022), MoleculeSTM (Liu et al., 2023a), MolCA-SMILES (Liu et al., 2023c), KV-PLM (Zeng et al., 2022), and Atomas (Zhang et al., 2025). Additionally, we compare with Uni-Mol (Zhou et al., 2023), which employs 3D molecular conformations for representation learning. A detailed baseline introduction can be found in Appendix D.